# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v1

@@ 2 A Dynamic Regional Model of Climate and the Economy with an ETS
This section introduces a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement and net zero emissions targets.
@@ 2.1 Emissions
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al. (2019). The amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. That is,
@@ 3.2 The Algorithm for the Noncooperative Model
This algorithm embodies the concept of market equilibrium. When there is a positive net quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices. Conversely, when there is a negative net quantity of traded emission permits, indicating excess supply, we lower the permit prices. This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears. Furthermore, this algorithm guarantees that each region obtains its optimal solution and reaches an equilibrium state. In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution. Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state.
@@ 4.1 SCC and MAC
It should be emphasized that our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al. (2018). Note that, for the global SCC, the marginal rate of substitution between global emissions and global capital (or consumption) is equivalent to the present value of future global climate damages from an additional unit of global emissions released in the current period. In this regard, it has been well understood in the literature that the global SCC can be used as the optimal global carbon tax. However, in Nordhaus (2017) and Ricke et al. (2018), their regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period. Such definition implies that their regional SCC is the regional contribution to the global SCC, i.e., the sum of their regional SCCs over all regions is the global SCC. This does not make sense in the context of carbon tax.
We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the OLNE. With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional net emissions is not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional net emissions is binding at zero. This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al. (2017) and Cai and Lontzek (2019). Similarly, the regional cooperative SCC in our definition is equal to the optimal regional carbon tax under cooperation, when the regional net emissions is not binding at zero and the exogenous emission permit price is zero (or there is no ETS). However, when the exogenous emission permit price is nonzero, the regional cooperative SCC is not equal to the optimal regional carbon tax under cooperation, because the ETS can be used as a means for transferring capital between regions to improve the global welfare under cooperation. For instance, a developed region can purchase a large amount of permits from a developing region even when its net emissions has been below its emission cap.
@@
When there is no ETS, it is well known that the optimal carbon tax is equal to the MAC. This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019). However, in the presence of the ETS, the equality does not hold as firms can purchase or sell emission permits instead of abatement.
@@ 4.2 Optimal Carbon Tax under the ETS
First of all, the relationship between the optimal net emission ($E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller. A region becomes a permit buyer if its optimal net emission is greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement. On the other hand, a region becomes a permit seller if its optimal net emission is lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$). Note that this is the optimal solution from the perspective of each region’s social planner.
@@
One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emission level. We assume that the social planner uses carbon taxation as the policy instrument to fix the inconsistency between the optimal net emissions of the regional social planner and the firms. Let us consider a representative firm in each region that produces emissions and faces the MAC function and the emission cap of the region. We assume that the representative firm is myopic and maximizes its profit in each period, thus it does not consider the dynamic optimization problem of the social planner. Given the total emissions, $E^{\mathrm{Total}}$, the representative firm’s profit maximization problem is equivalent to choosing its optimal level of net emissions to minimize its total cost associated with emissions (i.e. the abatement cost and revenue or expenditures from emission permit purchases or sales). [^4] It is worth mentioning that a region’s total emissions is proportional to its GDP, which, in turn, is determined by capital and the exogenous path of population size. Considering that the current level of capital is determined in the previous period, total emissions can be considered as exogenously given from the perspective of the single-period myopic firm.
@@
Figure 1 depicts the relationship between emissions, permit prices, the carbon tax, and the MAC for buyer and seller regions. In Figure 1, the MAC curve is a decreasing function of net emissions ($E$),  [^5] Substituting $E^{A}=E^{\mathrm{Total}}-E$ from equation (3) into the MAC, it follows that $\mathrm{MAC}_{i,t}=1,000\left(\frac{b_{1,i,t}b_{2,i}}{\sigma_{i,t}}\right)\left(\frac{E_{i,t}^{\mathrm{Total}}-E_{i,t}}{\sigma_{i,t}Q_{i,t}}\right)^{b_{2,i}-1}$. which reaches zero when net emissions equals total emissions ($E^{\mathrm{Total}}$). That is, the higher the emission abatement, the smaller the net emissions, resulting in a higher MAC. The market equilibrium price of emission permits is given at $m$. Then in the absence of a carbon tax, there is a corresponding cut-off level of net emissions, $E^{m},$ above which conducting the emission abatement is more cost-effective, and below which purchasing emission permits from the permit markets is cheaper.
@@
In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$. Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emission is smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits. That is, in the absence of a carbon tax, the social planner’s optimal net emission level, $E^{*}$, may not align with the optimal choice of firms within each region, $E^{m}$, regardless of whether the region is a permit buyer or seller. Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at $m$ to increase its profits. This is also verified in our simulation results later on.
@@
From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm leads to the following relationship between the carbon tax, the MAC, and the permit price:
$\tau=\mathrm{MAC}(E^{*})-m,$ (18)
when $E^{*}>0$. Importantly, it implies that for the representative firm to attain the optimal net emission level $E^{*},$ the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$. In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC of a region and the market price of emission permits. It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*},$ rather than $E^{m}$. When $E^{*}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
$\tau\geq\mathrm{MAC}(E^{*})-m.$ (19)
That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits. In particular, the regional $\mathrm{MAC}(E^{*})$ at $E^{*}=0$ could be smaller than the global permit price $m$, in which case the optimal carbon tax should be zero. Since the optimal carbon tax can always be chosen to be equal to the gap between the MAC and the permit price when the gap is positive, we have the following general formula for any $E^{*}\geq 0$:
$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\}.$ (20)
The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model. Therefore, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
@@ 5.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade or ETS currently in place, we use the emission commitments of the Paris Agreement and the regions’ net zero commitments to construct the baseline emission cap constraint, $\overline{E}_{i,t}$. We collect reports of the Intended Nationally Determined Contributions (INDCs) of 165 countries that were submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 (INDC, 2023). Although there is no international enforcement mechanism, most countries committed to achieving a self-defined mitigation target for 2030 (or 2025). We also obtain the target years to reach net zero emissions of different countries from Climate Action Tracker (2023). Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
As the first step, we obtain the level of emission targets by 2030 for each country. In their INDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level in 2015. Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in net emissions. For the countries that did not make a specific emissions reduction pledge, we assume that their carbon intensity reduction and emission reduction percentages are the same as those of the most populous country in that region.
@@ 5.4 Abatement Cost
Our estimation of the abatement cost parameters $b_{1,i}$, $b_{2,i}$, and $b_{3,i}$ relies on the simulation results under ten different levels of carbon taxes in Ueckerdt et al. (2019). For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$, Ueckerdt et al. (2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$. Since Ueckerdt et al. (2019) do not consider an ETS, according to the discussion in Section 4.2, we can assume that their carbon taxes $\tau_{t,j}^{U}$ are equal to the marginal abatement costs when net emissions are strictly positive. That is,
@@
where $E_{i,t,\mathrm{FFrun211}}^{U}$ is the regional emissions from the FFrun211 scenario with zero carbon tax (see Ueckerdt et al. (2019) for details). Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (24) can hold in an approximate manner for every scenario $j$ and time $t$.

### v2

@@ 2 A Dynamic Regional Model of Climate and the Economy with an ETS
This section introduces a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang, 1996; Nordhaus, 2010) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets. In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
@@ 2.1 Emissions
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al. (2019). We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the total industrial emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. [^5] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al., 2006; Bauer et al., 2012; Golosov et al., 2014; Baldwin et al., 2020). However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated. The emission control rate concept simplifies this. That is,
@@ 3.2 The Algorithm for the Noncooperative Model
This algorithm embodies the concept of market equilibrium. When there is a positive net quantity of traded emission permits in the market, indicating excess demand, we increase the permit prices. Conversely, when there is a negative net quantity of traded emission permits, indicating excess supply, we lower the permit prices. This mechanism ensures that the market reaches a balance between supply and demand, and thus the market clears. Furthermore, this algorithm guarantees that each region obtains its optimal solution and reaches an equilibrium state. In other words, no region has the incentive to deviate from the Nash equilibrium for the noncooperative model solution. Note that the equilibrium of the noncooperative model is an open-loop Nash equilibrium (OLNE), which provides a solution path over time depending on the initial state. The concept of the OLNE assumes the regions enter an agreement to commit to the solution path over time at the initial period. While this type of equilibrium concept might not be as satisfactory as the feedback Nash equilibrium concept, it has the computational advantages of solving open-loop versus feedback, particularly when the dimension of the state space is large and there are occasionally binding constraints as in our case.
@@ 4.1 SCC and MAC
Our concept of the regional SCC differs from the definition in Nordhaus (2017) and Ricke et al. (2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous. We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the OLNE. With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional net emissions is not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional net emissions is binding at zero. This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al. (2017) and Cai and Lontzek (2019). Moreover, we also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, but our social discount rates are endogenous.
Similarly, the regional cooperative SCC in our definition is equal to the optimal regional carbon tax under cooperation, when the regional net emissions is not binding at zero and the exogenous emission permit price is zero (or there is no ETS). However, when the exogenous emission permit price is nonzero, the regional cooperative SCC is not equal to the optimal regional carbon tax under cooperation, because the ETS can be used as a means for transferring capital between regions to improve the global welfare under cooperation. For instance, a developed region can purchase a large amount of permits from a developing region even when its net emissions has been below its emission cap. In addition, the regional cooperative SCC is not equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, because a developed region will abate more with a high carbon tax or cooperative SCC to improve the global welfare under cooperation.
@@ 4.2 Optimal Carbon Tax under the ETS
First of all, the relationship between the regional social planner’s optimal net emissions (denoted $E^{*}$) and emission cap ($\bar{E}$) determines whether a region becomes a permit buyer or seller. A region becomes a permit buyer if its optimal net emissions are greater than the emission cap, i.e., $E^{*}>\bar{E}.$ It follows that the emission quantity of ($E^{*}-\bar{E}$) becomes the optimal quantity of permits purchased and ($E^{\mathrm{Total}}-E^{*}$) is the optimal emission abatement. On the other hand, a region becomes a permit seller if its optimal net emissions are lower than its emission cap, i.e., $E^{*}<\bar{E}$, with the optimal quantity of permits sold captured by ($\bar{E}-E^{*}$). Note that this is the optimal solution from the perspective of each region’s social planner.
@@
One issue arises when the social planner of each region cannot internally force its firms—the actual emission producers in the economy—to obtain the social planner’s optimal net emissions level. We assume that the social planner uses carbon taxation as the policy instrument to fix the inconsistency between the optimal net emissions of the regional social planner and the firms. Let us consider a representative firm in each region that produces emissions and faces the MAC function and the emission cap of the region. We assume that the representative firm is myopic and maximizes its profit in each period, thus it does not consider the dynamic optimization problem of the social planner. Given the total emissions, $E^{\mathrm{Total}}$, the representative firm’s profit maximization problem is equivalent to choosing its optimal level of net emissions to minimize its total cost associated with emissions (i.e. the abatement cost and revenue or expenditures from emission permit purchases or sales). [^6] It is worth mentioning that a region’s total emissions is proportional to its GDP, which, in turn, is determined by capital and the exogenous path of population size. Considering that the current level of capital is determined in the previous period, total emissions can be considered as exogenously given from the perspective of the single-period myopic firm.
@@
Figure 1 depicts the relationship between emissions, permit price, carbon tax, and the MAC for buyer and seller regions. In Figure 1, the MAC curve is a decreasing function of net emissions ($E$),  [^7] Substituting $E^{A}=E^{\mathrm{Total}}-E$ from equation (3) into the MAC, it follows that $\mathrm{MAC}_{i,t}=1,000\left(\frac{b_{1,i,t}b_{2,i}}{\sigma_{i,t}}\right)\left(\frac{E_{i,t}^{\mathrm{Total}}-E_{i,t}}{\sigma_{i,t}Q_{i,t}}\right)^{b_{2,i}-1}$. which reach zero when net emissions equal total emissions ($E^{\mathrm{Total}}$). That is, the higher the emission abatement, the smaller the net emissions, resulting in a higher MAC. The market equilibrium price of emission permits is given at $m$. Then in the absence of a carbon tax, there is a corresponding cut-off level of net emissions, $E^{m},$ above which conducting the emission abatement is more cost-effective, and below which purchasing emission permits from the permit markets is cheaper.
@@
In the case of a permit buyer region (the left panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ then its MAC is larger than the permit price $m$, so the firm has an incentive to purchase a quantity of $(E^{m}-\bar{E})$ emission permits and reduce its abatement effort until its net emissions become $E^{m}$. Conversely, in a permit seller region (the right panel of Figure 1), if there is no carbon tax, when the firm’s net emissions are smaller than $E^{m},$ it is more profitable for the firm to reduce its abatement effort until its net emissions reach $E^{m}$ and then sell a quantity of $(\bar{E}-E^{m})$ emission permits. That is, in the absence of a carbon tax, the social planner’s optimal net emissions, $E^{*}$, may not align with the optimal choice of the firm within a region, $E^{m}$, regardless of whether the region is a permit buyer or seller. For example, when $\bar{E}$ is very large for every region, the permit price $m$ is zero, so $E^{m}=E^{\mathrm{Total}}$, but it is well known in the literature that $E^{*}$ is often strictly smaller than $E^{\mathrm{Total}}$ (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023). Note that it is impossible for the firm’s net emissions to be larger than $E^{m}$ (i.e., its MAC to be smaller than $m$), because the firm as a buyer is not willing to purchase emission permits at a price $m$ that exceeds its $\mathrm{MAC}$ (it would instead reduce its costs via more abatement), and the firm as a seller would prefer to reduce its net emissions so it can sell more permits at the price $m$ to increase its profits. This is also verified in our simulation results later on.
@@
From the Karush-Kuhn-Tucker conditions of (17), the solution to the optimal net emissions of a firm (denoted $E^{*,\mathrm{Firm}}$) leads to the following relationship between the carbon tax, the MAC, and the permit price:
$\tau=\mathrm{MAC}(E^{*,\mathrm{Firm}})-m,$ (18)
when $E^{*,\mathrm{Firm}}>0$. Importantly, it implies that for the representative firm to attain the regional social planner’s optimal net emissions $E^{*}$ (i.e., $E^{*,\mathrm{Firm}}=E^{*}$), the social planner can impose an internal carbon tax at $\mathrm{MAC}(E^{*})-m$, which is also equal to the regional SCC in our definition when $E^{*}>0$. In other words, the optimal regional carbon tax is endogenously determined by the gap between the MAC at the regional social planner’s optimal net emissions and the market price of emission permits. It aligns with the graphical illustration in Figure 1, capturing the idea that the social planner of a region can choose the level of carbon tax such that the firm can internalize the regional SCC and achieve its net emissions level of $E^{*,\mathrm{Firm}}=E^{*}$, rather than $E^{m}$. When $E^{*,\mathrm{Firm}}=0$, the Karush-Kuhn-Tucker conditions of (17) shows that
$\tau\geq\mathrm{MAC}(E^{*,\mathrm{Firm}})-m.$ (19)
That is, the carbon tax could be larger than the gap between the MAC of a region and the market price of emission permits. In particular, the regional $\mathrm{MAC}(E^{*,\mathrm{Firm}})$ at $E^{*,\mathrm{Firm}}=0$ could be smaller than the global permit price $m$. However, when $E^{*}=0$, a nonnegative carbon tax larger than $\mathrm{MAC}(E^{*})-m$ will not make $E^{*,\mathrm{Firm}}$ to be negative, so the optimal carbon tax should be zero or the gap $\mathrm{MAC}(E^{*})-m$ if it is positive.
Therefore, for any regional social planner’s optimal net emissions $E^{*}\geq 0$, we have the following general formula to choose the optimal carbon tax:
$\tau=\max\{0,\ \mathrm{MAC}(E^{*})-m\},$ (20)
such that the firm’s optimal net emissions are consistent with the social planner’s (i.e., $E^{*,\mathrm{Firm}}=E^{*}$). The simple relationship between the optimal carbon tax, the MAC, and the permit price is also verified in our numerical solutions of the dynamic regional model. Thus, when we compare the regional SCC to the regional optimal carbon tax in our numerical results, we will compare it to the gap of the regional MAC and the permit price.
4.3 Optimal Carbon Tax under no ETS
For the case without the ETS, the cost minimization problem of the representative firm facing the emission cap regulation becomes:
[…]
From the Karush-Kuhn-Tucker conditions of (21), when the firm’s optimal solution $E^{*,\mathrm{Firm}}=E^{*}$ is not binding at its lower or upper bound, we have the well known equality:
$\tau=\mathrm{MAC}(E^{*}).$ (22)
That is, the optimal carbon tax should be chosen to be equal to the MAC at the social planner’s optimal emissions $E^{*}$ if it is not binding. This equality has been often used in the literature (e.g., Nordhaus, 2014; Cai et al., 2017; Cai and Lontzek, 2019). We can also see that when $\overline{E}$ is very large or the cap does not exist, a nonzero $E^{*}$ will not be binding, so the equality holds, then the MAC is nonzero. In the case with the ETS and very large caps for every region, which is equivalent to the case without the ETS nor the caps, the permit price should be zero, and so our formula (20) is consistent with (22). However, the assumption that the MAC is equal to the permit price, which is often used in the literature (e.g., Keohane, 2009; Nordhaus, 2010), does not hold in this case, unless $E^{*}=E^{\mathrm{Total}}$ (i.e., the social planner’s optimal decision is to have zero abatement), which usually does not appear in those non-business-as-usual solutions in the literature (see, e.g., Nordhaus and Yang, 1996; Cai et al., 2023).
@@ 5.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade or ETS currently in place, we use the emission commitments of the Paris Agreement and the regions’ net zero commitments to construct the baseline emission cap constraint, $\overline{E}_{i,t}$. We collect reports of the current Nationally Determined Contributions (NDCs) of 165 countries that were first submitted at the United Nations Framework Convention on Climate Change (UNFCCC) Conference of the Parties (COP 21) in Paris in December 2015 and later updated (INDC, 2023; NDC, 2024). Although there is no international enforcement mechanism, most countries committed to achieving a self-defined mitigation target for 2030 (or 2025). We also obtain the target years to reach net zero emissions of different countries from Climate Action Tracker (2023). Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
As the first step, we obtain the level of emission targets for each country. In their NDC reports, most countries expressed targets as a specific percentage reduction in emissions by 2030 (or, in some cases, 2025) compared to their Business As Usual (BAU) emission level at some base year. Some countries, including China, Chile, Malaysia, Singapore, and Tunisia specified their targets as a percentage reduction in carbon intensity instead of a percentage reduction in net emissions. For the countries that did not make a specific emissions reduction pledge, we assume that their carbon intensity reduction and emission reduction percentages are the same as those of the most populous country in that region.
@@ 5.4 Abatement Cost
Our estimation of the abatement cost parameters $b_{1,i}$, $b_{2,i}$, and $b_{3,i}$ relies on the simulation results under ten different levels of carbon taxes in Ueckerdt et al. (2019). For each scenario $j$ with associated carbon taxes $\tau_{t,j}^{U}$ for every region $i$, Ueckerdt et al. (2019) report the projected regional emissions net of abatement ($E_{i,t,j}^{U}$), for region $i$ at time $t$. Since Ueckerdt et al. (2019) do not consider an ETS, according to the discussion in Section 4.2, we can assume that their carbon taxes $\tau_{t,j}^{U}$ are equal to the marginal abatement costs when net emissions are strictly positive. That is,
@@
where $E_{i,t,\mathrm{FFrun211}}^{U}$ is the regional emissions from the FFrun211 scenario with zero carbon tax (see Ueckerdt et al. (2019) for details). Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (26) can hold in an approximate manner for every scenario $j$ and time $t$. Appendix A.8.6 also reports a sensitivity analysis on the estimate of the abatement cost.